OpenStudy (anonymous):

6. Starting from 1.5 miles away, a car drives toward a speed checkpoint and then passes it. The car travels at a constant rate of 53 miles per hour. The distance of the car from the checkpoint is given by d = |1.5 – 53t|. At what times is the car 0.1 miles from the checkpoint? Calculate your answer in seconds. (1 point) 95.1s and 108.7s 10.2s and 101.9s 108.7s and 10.2s 95.1s and 10.2s

OpenStudy (anonymous):

strange way to ask: "solve \(|1.5-53t|=0.1\) for \(t\)

OpenStudy (anonymous):

how do u slove?

OpenStudy (anonymous):

|1.5−53t| = 0.1 implies 1.5 - 53t = 0.1 and 53t - 1.5 = 0.1

OpenStudy (anonymous):

answer?

OpenStudy (anonymous):

you just solve for t in each equation to get each answer. your values of t will be in hours, so you have to convert them into seconds

OpenStudy (anonymous):

3180

OpenStudy (anonymous):

i'll do the first one for you 1.5 - 53t = 0.1 1.5 = 0.1 + 53t 1.4 = 53t 0.026415 = t 0.026415 * 60 min/h * 60 s/min = 95.094 seconds or about 95.1 seconds

OpenStudy (anonymous):

now i c wat u do thanks